OpenStudy (loser66):

For which integers n is phi (n) = n/2 Please, help

OpenStudy (loser66):

phi (n) is an integer, hence n must be an even number to get n/2 = integer hence \(n = 2^a k\) where k is odd.

OpenStudy (ikram002p):

i think that goes 2^k numbers we can prove that

OpenStudy (ikram002p):

u are on the right track lol

OpenStudy (loser66):

\(\phi (n) = \phi (2^a) *\phi (k) = 2^{a-1}k\), then?

OpenStudy (ikram002p):

well if there is k odd then there is some odd integer lets say i<=k, such that gcd(i,n) is not 1

OpenStudy (ikram002p):

so k need to be even thus k=2h and again we would have another integer added to n/2 so we kinda reach k can be of the form 2^g only.

OpenStudy (loser66):

no way!! n is even, i < = k is odd, how (i, n ) is not 1?

OpenStudy (loser66):

k is not even, because if we set n = 2^a k, then k must be odd. (all even goes to 2^)

OpenStudy (ikram002p):

=) let n=6 gcd(6,3)=3 for example.

OpenStudy (loser66):

oh yeah, my stupidity. :)

OpenStudy (ikram002p):

nope its ok, i wanted u to see the contradiction to find out that k can be of the form 2^g or something.

OpenStudy (ikram002p):

so lets try to prove this shall we ? phi(n)=n/2 iff n=2^k

OpenStudy (loser66):

show me, please

OpenStudy (ikram002p):

<--- when n=2^k then for all even numbers 2m, gcd(2m,n) is not 1. (maybe 2 or 2^i for i<k), and we have from 1 to n, n/2 even numbers.

OpenStudy (loser66):

is it not that we need prove --> if n = 2^k, then phi (n) = n/2 <--, if phi (n) = n/2, then n=2^k

OpenStudy (ikram002p):

--->ph(n)=n/2 then n is even , let \( n= 2^{a_1}p_1^{a_2}p_2^{a_3}....\) so for all evens 2m gcd(2m, n) is not 1, means phi(n)>=n/2 as we have other prime factors for n but not 2, so the only case we have is when n have only prime factors of 2.

OpenStudy (ikram002p):

oh! @Loser66 do i make sense to u or i just confused u ?

OpenStudy (loser66):

I don't get why \(gcd (2m,n)\neq 1, ~then~~\phi(n)\geq n/2\)

OpenStudy (ikram002p):

ok so phi(n)=n/2 means n even right ?

OpenStudy (loser66):

yes

OpenStudy (ikram002p):

so for even numbers 2m from m=1 to m=n/2 gcd(2m,n) is not 1 (it can be 2 or 2i) right ?

OpenStudy (ikram002p):

oh sorry that inequality should be phi(n) <=n/2 pardon.

OpenStudy (loser66):

it's ok, got you

OpenStudy (ikram002p):

so now at most we have n/2 of numbers are relatively prime with n, assume n=2^ak and k is odd factor, then phi(n)<n/2.

OpenStudy (loser66):

ok

OpenStudy (loser66):

I got it, friend. Thank you so much:)

OpenStudy (ikram002p):

you are welcome <3